OpenStudy (anonymous):

A line has a slope of 2 and contains the point (3,5). What is the equation of the line?

OpenStudy (anonymous):

use this formula y-y1=m(x-x1) here x1=3, y=5 and m=slope = 2 do it

OpenStudy (anonymous):

sorry its y1 = 5

OpenStudy (anonymous):

y-5 = 2(x-3) which becomes y-5= 2x -6. add the 5 to both sides and you get y=2x-1

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

thanks for clearing that up. As soon as you started typing how to do the equation, I remembered how to do it.. haha

OpenStudy (anonymous):

haha thats nice